OpenStudy (like1234):

could someone show me how to solve this (5 • 8 – 12)(36 ÷ 9 • 4)

OpenStudy (perl):

use order of operations. PEMDAS

OpenStudy (anonymous):

you got to use order of operation perenthisis multiplication division addition and last but not least subtraction if you have that in your problem

OpenStudy (anonymous):

5 * 8= 40 40 - 12= 28 36/9= 4 4 * 4= 16 finally: 16 * 28= 448

OpenStudy (like1234):

thank you so much! :)

OpenStudy (anonymous):

No prob :)